Not from book!
A car machanic needs to mix together two solutions of anti-freeze. One solution is 100% pure anti-freeze and the other is 30% anti-freeze. The machanic would like to have 2.5 gallons of a mixture that is 44% anti-freeze. How many gallons of each solution will the car machanic need?

Answer by Nate(3500)

(1x + 0.3(2.5 - x))/2.5 = 0.44
x + 0.3(2.5 - x) = 1.1
x + 0.75 - 0.3x = 1.1
0.7x + 0.75 = 1.1
0.7x = 0.35
x = 0.5 gallons 100%
2.5 - x = 2 gallons 30%